The internal energy increases by 25 J as a gas-filled piston expands 0.30 m as the piston is pushed outward by a 70 N force. What is the heat added/removed from the system?


1
Expert's answer
2021-03-16T11:35:53-0400

By the first law of thermodynamics, we have:


"\Delta U=Q-W,"

here, "\Delta U" is the change in the internal energy of the gas, "Q" is the heat added or removed from the gas, "W" is the work done by the gas (since the gas volume expands, the work done by the gas has negative sign).

Then, we can find "Q":


"Q=\Delta U+W=\Delta U+Fd,""Q=25\ J+70\ N\cdot0.3\ m=46\ J."

Since "Q" is positive, heat is added to the system. Therefore, 46 J of heat added to the system.